Find the electric flux θ∈ if the electric field E with a value of 5.78 μC is contained in a conical container with a radius of 1.8 m and a height of 4.56 m

 



1
Expert's answer
2022-05-22T14:39:03-0400

The Gauss' law says

"\theta=\oint E dA=\frac{q}{\epsilon_0}"

"\theta=\frac{5.78*10^{-6}}{8.85*10^{-12}}=6.53*10^5\:\rm N\cdot m^2/C"